OpenStudy (anonymous):

What is the value of l in the following formula when w = 14 cm and P = 80 cm? P = 2l + 2w

OpenStudy (across):

What have you tried?

OpenStudy (anonymous):

I am not sure how to do this. It would be better if someone was able to walk through with me on this.

OpenStudy (anonymous):

P=2L+2W 80=2L+2(14) 80=2L+28 80-28=2L+28-28 52=2L 52/2=2L/2 26=L

OpenStudy (anonymous):

Oh. That really is not that hard. I had the idea correct, I just didnt know where the numbers were supposed to go. THanks!

OpenStudy (across):

You are told to solve for \(l\) in\[P=2l+2w.\]Moreover, you are told that \(P=80\) and \(w=14\). If you substitute these values into the above equation, you get\[80=2l+2\times14.\]Can you simplify this?

OpenStudy (anonymous):

You're welcome!